# MrSteve: Instruction-Following Agents in Minecraft with What-Where-When Memory

### v4

@@ Appendix I MrSteve in Randomly Generated Maps
We report success rates of DEPS framework with Steve-1 and MrSteve in Table 9. We use Qwen2.5-72B (Team, 2024) as LLM in the high-level planner. DEPS with MrSteve shows comparable performance or outperforms compared to DEPS with Steve-1. Especially, there is a huge performance gap between the two low-level controllers in the bed task, which requires killing three sheeps. We observe that when the agent hit a sheep, it runs away from the agent and the agent chases it until the agent gets wool items. After this, Steve-1 easily forgets the place where other sheeps exist and try to find sheep. However, MrSteve avoids this redundant exploration utilizing the episodic memory. Hence, this results highlight the importance of episodic memory for low-level controllers in LLM-augmented hierarchical agent frameworks.
@@
In this section, we investigate how MrSteve and its memory variants perform in ABA-Sparse tasks in Section 4.2 when memory capacity is limited. We evaluated each model with different memory capacity ranging from 0.1K to 12K which is the maximum episode length. In all tasks, we observe that the performance of MrSteve-FM decreases as the memory capacity gets small. This is because FIFO Memory in MrSteve-FM losts relevant frames in first task A while solving task B. While MrSteve showed robust performance to constrained memory capacities across tasks, MrSteve-PM, and MrSteve-EM showed degraded performances in Beef-Log-Beef and Beef-Leaves-Beef tasks when memory capacity is 0.1K. This indicates the robustness of PEM to memory capacities in ABA-Sparse tasks.

### v5

@@ Appendix I MrSteve in Randomly Generated Maps
We report success rates of DEPS framework with Steve-1 and MrSteve in Table 9. We use Qwen2.5-72B (Team, 2024) as LLM in the high-level planner. DEPS with MrSteve shows comparable performance or outperforms compared to DEPS with Steve-1. Especially, there is a huge performance gap between the two low-level controllers in the bed task, which requires killing three sheeps. We observe that when the agent hit a sheep, it runs away from the agent and the agent chases it until the agent gets wool items. After this, Steve-1 easily forgets the place where other sheeps exist and tries to find sheep. However, MrSteve avoids this redundant exploration by utilizing the episodic memory. Hence, these results highlight the importance of episodic memory for low-level controllers in LLM-augmented hierarchical agent frameworks.
@@
In this section, we investigate how MrSteve and its memory variants perform in ABA-Sparse tasks in Section 4.2 when memory capacity is limited. We evaluated each model with different memory capacities ranging from 0.1K to 12K, which is the maximum episode length. In all tasks, we observe that the performance of MrSteve-FM decreases as the memory capacity gets small. This is because FIFO Memory in MrSteve-FM losts relevant frames in first task A while solving task B. While MrSteve showed robust performance to constrained memory capacities across tasks, MrSteve-PM, and MrSteve-EM showed degraded performances in Beef-Log-Beef and Beef-Leaves-Beef tasks when memory capacity is 0.1K. This indicates the robustness of PEM to memory capacities in ABA-Sparse tasks.